Among 157 African American men, the mean systolic blood pressure was 146 mm Hg with a standard deviation of 27. We wish to know if based on these data, we may conclude that the mean systolic blood pressure for a population of African American is greater than 140. Use α=0.01.

1
Expert's answer
2022-02-06T16:30:10-0500

"n=157\\\bar x=146\\s=27"

The hypothesis tested are,

"H_0:\mu=140\\vs\\H_1:\mu\gt140"

The test statistic is given as,

"t={\bar x-\mu\over{s\over\sqrt{n}}}={146-140\over{27\over\sqrt{157}}}={6\over2.155}=2.7844"

The critical value is the t distribution table value at "\alpha=0.01"with "n-1=157-1=156" degrees of freedom given as, "t_{0.01,156}= 2.350489"

The null hypothesis is rejected if, "t\gt t_{0.01,156}"

For this case, "t=2.7844\gt t_{0.01,156}=2.350489" thus, we reject the null hypothesis and conclude that there is sufficient evidence to show that the mean systolic blood pressure for a population of African American is greater than 140 at 1% significance level.